# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 6.2 Wasserstein Distance Guarantees
Table 1: Summary of the bounds on the Wasserstein distance between a distribution $p\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ and $q\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ obtained by pushing forward $p$ through the most common SNN operation types. Here, ${\mathcal{L}}_{\sigma}$ and ${\mathcal{L}}_{\sigma\mid{\mathcal{R}}_{i}}$ are, respectively, the global Lipschitz constant and the local Lipschitz constant in region ${\mathcal{R}}_{i}$ of the piecewise continuous activation function $\sigma$, [^11] For ReLU, ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}=1$ for regions ${\mathcal{R}}_{i}$ overlapping with the positive half plain, and zero else. For tanh, ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}\in(0,1]$. while ${\mathbb{W}}_{2\mid{\mathcal{R}}_{i}}^{2}$ is as defined in Proposition 8. The propositions and their proofs are provided in Appendix A.3.
@@ 6.3 Approximation for Sets of Input Points
We now consider the case where ${\mathcal{X}}=\{{{x}}_{1},...,{{x}}_{D}\}$, that is, a finite set of input points, and extend our approach to compute a Gaussian mixture approximation of $p_{nn({\mathcal{X}})}$. We first note that $p_{nn({\mathcal{X}})}$ can be equivalently represented by extending Eqn. (4) as follows, where in the equation below ${\mathrm{vec}}({\mathcal{X}})=({{x}}_{1}^{T},\ldots,{{x}}_{D}^{T})^{T}$ is the vectorization of ${\mathcal{X}}$
$p_{nn({\mathcal{X}}),1}={\mathbb{E}}_{{{z}}_{0}\sim\delta_{{\mathrm{vec}}({\mathcal{X}})}}[p_{L_{{{w}}_{0}}^{D}({{z}}_{0})}],$ (21)
$p_{nn({\mathcal{X}}),k+1}={\mathbb{E}}_{{{z}}_{k}\sim p_{nn({\mathcal{X}}),k}}[p_{L^{D}_{{{w}}_{k}}({\sigma}({{z}}_{k}))}],\qquad k\in\{1,...,K\},$
$p_{nn({\mathcal{X}})}=p_{nn({\mathcal{X}}),K+1},$
@@
where $L_{{{w}}_{k}}^{D}:{\mathbb{R}}^{D\cdot n_{k}}\rightarrow{\mathbb{R}}^{D\cdot n_{k}}$ is the stacking of $D$ times $L_{k}$, i.e., $L_{{{w}}_{k}}^{D}(({{z}}_{1}^{T},\ldots,{{z}}_{D}^{T})^{T})=(L_{{{w}}_{k}}({{z}}_{1})^{T},\ldots,\newline L_{{{w}}_{k}}({{z}}_{D})^{T})^{T}$, and $p_{L^{D}_{{{w}}}({{z}})}=L^{D}_{{{w}}_{k}}({{z}})\#p_{{{w}}_{k}}$. That is, $p_{nn({\mathcal{X}})}$ is computed by stacking $D$ times the neural network for the inputs in ${\mathcal{X}}$. Following the same steps as in the previous section, $p_{nn({\mathcal{X}})}$ can then be approximated by the Gaussian Mixture distribution $q_{nn({\mathcal{X}})}$, for $k\in\{1,\ldots,K\}$ defined as
@@
$q_{nn({\mathcal{X}}),1}$ $={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{\mathrm{vec}}({\mathcal{X}})}}[p_{L^{D}_{{{w}}_{0}}({{{z}}}_{0})}]$ (Initialization and Linear operation) (22a)
$d_{nn({\mathcal{X}}),k}$ $={\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({\mathcal{X}}),k})$ (Compression and Signature operations) (22b)
@@
$q_{nn({\mathcal{X}}),k+1}$ $={\mathbb{E}}_{{{{z}}}_{k}\sim d_{nn({\mathcal{X}}),k}}[p_{L^{D}_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}]$ (Activation and Linear operations) (22c)
$q_{nn({\mathcal{X}})}$ $=q_{nn({\mathcal{X}}),K+1}$ (Output) (22d)
@@
where ${\mathcal{C}}_{k}\subset{\mathbb{R}}^{D\cdot n_{k}}$ are the signature locations at layer $k$, and $M\in\mathbb{N}$ is the size of compressed mixtures. A corollary of Theorem 13 is the following proposition that bounds the approximation error of $q_{nn({\mathcal{X}})}$.
@@
Let $p_{nn({\mathcal{X}})}$ be the joint distribution of a SNN with $K$ hidden layers over a finite set of input points ${\mathcal{X}}=\{{{x}}_{i}\}_{i=1}^{D}\subset{\mathbb{R}}^{n_{0}}$ and $q_{nn({\mathcal{X}})}$ be a Gaussian mixture distribution built according to Eqn. (22). For $k\in\{1,\ldots,K\}$ iteratively define $\hat{{\mathbb{W}}}_{2}$ as in Eqn. (14) with $\hat{\mathcal{S}}_{2,{{w}},k}$, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ and $\hat{{\mathbb{W}}}_{2,{R},k}$ such that
$D^{\frac{1}{2}}{\mathbb{E}}_{{{{W}}}_{k}\sim p_{w_{k}}}[\|{{{W}}}_{k}\|^{2}]^{\frac{1}{2}}$ $\leq\hat{\mathcal{S}}_{2,{{w}},k},$ (23)
${\mathbb{W}}_{2}({\sigma}\#{R}_{M}(q_{nn({\mathcal{X}}),k}),{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({\mathcal{X}}),k}))$ $\leq\hat{{\mathbb{W}}}_{2,{\Delta},k},$ (24)
${\mathbb{W}}_{2}(q_{nn({\mathcal{X}}),k},{R}_{M}(q_{nn({\mathcal{X}}),k}))$ $\leq\hat{{\mathbb{W}}}_{2,{R},k}.$ (25)
Then, it holds that
In the above proposition, we can set $\hat{\mathcal{S}}_{2,{{w}},k}$ as in the single-point case, scaled by the square root of the number of input points. Similar to the single point-case, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ can be computed as in Eqn. (12) using Corollaries 10 and 27, and $\hat{{\mathbb{W}}}_{2,{R},k}$ can be taken as the ${\mathbb{MW}}_{2}$ distance between GMMs $q_{nn({\mathcal{X}}),k}$ and ${R}_{M}(q_{nn({\mathcal{X}}),k})$. Note that the effect of multiple input points on the error introduced by $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ and $\hat{{\mathbb{W}}}_{2,{R},k}$ depends on the specific choice of points in ${\mathcal{X}}$. For points in ${\mathcal{X}}$ where $q_{nn({\mathcal{X}}),k}$ is highly correlated, the error will be similar to that of a single-input case. However, if there is little correlation, the error bound will increase linearly with the number of points.
6.4 Convergence Analysis
In Theorem 13 and Proposition 16 we derived error bounds on the distance between $q_{nn({\mathcal{X}})}$ and $p_{nn({\mathcal{X}})}$. In the following theorem, we prove that if one allows for the size of $q_{nn({{x}})}$ to be arbitrarily large, then the error converges to zero.
Theorem 17
Let $p_{nn({\mathcal{X}})}$ be the output distribution of a SNN with $K$ hidden layers for a finite set of inputs ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$, $q_{nn({\mathcal{X}})}$ be a Gaussian mixture distribution built according to Eqn. (22), and $\hat{{\mathbb{W}}}_{2}$ be the upper bound on ${\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})$ defined according to Eqn. (14). Then, for any $\epsilon>0$, there exist sets of signature locations ${\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{K}$ of finite sizes, and a compression size $M$, such that $\hat{{\mathbb{W}}}_{2}\leq\epsilon$.
[…]
Remark 18
While in the proof of Theorem 17 we use uniform grids for each component of $q_{nn({\mathcal{X}}),k}$, in practice, adaptive, non-uniform gridding can guarantee closeness in ${\mathbb{W}}_{2}$-distance with the same precision for smaller grid sizes.
7 Algorithmic Framework
In this section, the theoretical methodology to solve Problem 1 developed in the previous sections is translated into an algorithmic framework. First, following Eqn. (22), we present an algorithm to construct the Gaussian mixture approximation of a SNN with error bounds in terms of the 2-Wasserstein distance, including details on the compression step. Then, we rely on the fact that the error bound resulting from our approach is piecewise differentiable with respect to the parameters of the SNN to address the complementary problem highlighted in Remark 4. This problem involves deriving an algorithm to optimize the parameters of a SNN such that the SNN approximates a given GMM.
input : $p_{nn({\mathcal{X}})}$
output : $q_{nn({\mathcal{X}})}$; $\hat{{\mathbb{W}}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})$
begin
      1 Initialize $q_{nn({\mathcal{X}}),1}$ as in Eqn. (22a) and $\hat{{\mathbb{W}}}_{2,1}$ as in Eqn. (14a)
       for $k\in\{1,\ldots,K\}$ do
            2 Compress $q_{nn({\mathcal{X}}),k}$ using Algorithm 3 [Eqn. (22b)]
            3 Compute $\hat{{\mathbb{W}}}_{2,{R},k}$ using the ${\mathbb{MW}}_{2}$ distance [Eqn. (25)]
            4 Get signature $d_{nn({\mathcal{X}}),k}$ of ${R}_{M}(q_{nn({\mathcal{X}}),k})$ and $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ via Alg. 2 [Eqns. (22b) & (24)]
            5 Construct $q_{nn({\mathcal{X}}),k+1}$ as in Eqn. (22c)
            6 Compute $\hat{\mathcal{S}}_{2,{{w}},k}$ using Lemma 14 [Eqn. (23)]
            7 Compute $\hat{{\mathbb{W}}}_{2,k+1}$ as in Eqn. (14c)
      8Set $q_{nn({\mathcal{X}})}=q_{nn({\mathcal{X}}),K+1}$ and $\hat{{\mathbb{W}}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})=\hat{{\mathbb{W}}}_{2,K+1}$ [Eqns. (22d) & (14c)]
Algorithm 1 Gaussian Mixture Approximation of a SNN
7.1 Construction of the GMM Approximation of a SNN
The procedure to build a Gaussian mixture approximation $q_{nn({\mathcal{X}})}$ of $p_{nn({\mathcal{X}})}$, i.e. the output distribution of a SNN at a set of input points ${\mathcal{X}}$ as in Eqn (22), is summarized in Algorithm 1. The algorithm consists of a forward pass over the layers of the SNN. First, the Gaussian output distribution after the first linear layer is constructed (line 1). Then, for each layer $k>0$, $q_{nn({\mathcal{X}}),k}$ is compressed to a mixture of size $M$ (line 2); the signature of $R_{M}(q_{nn({\mathcal{X}})})$ is computed (line 4); and the signature is passed through the activation and linear layer to obtain $q_{nn({\mathcal{X}}),k+1}$ (line 5). Parallel to this, following Proposition 16, error bounds on the 2-Wasserstein distance are computed for the compression operation (line 3), signature and activation operation (line 4), and linear operation (line 6), which are then composed into a bound on ${\mathbb{W}}_{2}(p_{nn({\mathcal{X}}),k+1},q_{nn({\mathcal{X}}),k+1})$ that is propagated to the next layer (line 8).
Signature Operation
The signature operation on $q_{nn({\mathcal{X}}),k}$ in line 4 of Algorithm 1 is performed according to the steps in Algorithm 2. That is, the signature of $q_{nn({\mathcal{X}}),k}$ is taken as the mixture of the signatures of the components of $q_{nn({\mathcal{X}}),k}$ (lines 1-7). For each Gaussian component of the mixture, to ensure an analytically exact solution for the Wasserstein distance resulting from the signature, we position its signature locations on a grid that aligns with the covariance matrix’s basis vectors, i.e., the locations satisfy the constraint as in Eqn. (10) of Corollary 10. In particular, the grid of signatures is constructed so that the Wasserstein distance from the signature operation is minimized. According to the following proposition, the problem of finding this optimal grid can be solved in the transformed space, in which the Gaussian is independent over its dimensions.
Proposition 19
@@ 7.1 Construction of the GMM Approximation of a SNN
For $\mathcal{N}({m},\Sigma)\in{\mathcal{P}}({\mathbb{R}}^{n})$, let matrix ${{T}}\in{\mathbb{R}}^{n\times n}$ be such that ${{T}}=\mathrm{diag}({{\lambda}})^{-\frac{1}{2}}{{V}}^{T}$ where $\mathrm{diag}({{\lambda}})={{V}}^{T}\Sigma{{V}}$ is a diagonal matrix whose entries are the eigenvalues of $\Sigma$, and ${{V}}$ is the corresponding orthogonal (eigenvector) matrix. Then,
@@
$\underset{\begin{subarray}{c}\{\bar{{\mathcal{C}}}^{1},\ldots,\bar{{\mathcal{C}}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|\bar{{\mathcal{C}}}^{j}|=N\end{subarray}}{\mathrm{argmin}}{\mathbb{W}}_{2}({\Delta}_{\mathrm{Post}(\bar{{\mathcal{C}}^{1}}\times\ldots\times\bar{{\mathcal{C}}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))=\{\bar{{\mathcal{C}}}(N_{1}^{*}),\ldots,\bar{{\mathcal{C}}}(N_{n}^{*})\},$

### v2

@@ 6.2 Wasserstein Distance Guarantees
Table 1: Summary of the bounds on the $2$-Wasserstein distance between two distributions $p,q\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, obtained by pushing forward $p$ and $q$ through common SNN operation types. Here, ${\mathcal{L}}_{L^{{w}}}$ denotes the global Lipschitz constant of a Lipschitz function $L^{{w}}$, parametrized by ${{w}}\in{\mathbb{R}}^{d}$ and distributed according to $p_{{w}}\in{\mathcal{P}}({\mathbb{R}}^{d})$, and ${\mathcal{L}}_{\sigma}$ and ${\mathcal{L}}_{\sigma\mid{\mathcal{R}}_{i}}$ denote, respectively, the global Lipschitz constant and the Lipschitz constant over region ${\mathcal{R}}_{i}$ of a Lipschitz function $\sigma$. [^13] For ReLU, ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}=1$ for regions ${\mathcal{R}}_{i}$ overlapping with the positive half-space, and 0 else. For tanh, ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}\in(0,1]$. Proofs of these results are provided in Appendix B.2.
@@ 6.3 Approximation for Sets of Input Points
We now consider the case where ${\mathcal{X}}=\{{{x}}_{1},...,{{x}}_{D}\}$, that is, a finite set of input points, and extend our approach to a Gaussian mixture approximation of $p_{nn({\mathcal{X}})}$. We first note that $p_{nn({\mathcal{X}})}$ can be equivalently represented by extending Eqn (4) as follows, where in the equation below ${\mathrm{vec}}({\mathcal{X}})=({{x}}_{1}^{T},\ldots,{{x}}_{D}^{T})^{T}$ is the vectorization of $\mathcal{X}$},
$p_{nn({\mathcal{X}}),1}={\mathbb{E}}_{{{z}}_{0}\sim\delta_{{\mathrm{vec}}({\mathcal{X}})}}\left[\bm{L}^{{{w}}_{0}}({{z}}_{0})\#p_{{{w}}_{0}}\right],$ (24a)
$p_{nn({\mathcal{X}}),k+1}={\mathbb{E}}_{{{z}}_{k}\sim p_{nn({\mathcal{X}}),k}}\left[\bm{L}^{{{w}}_{k}}({\sigma}({{z}}_{k}))\#p_{{{w}}_{k}}\right],\qquad k\in\{1,\ldots,K\},$ (24b)
$p_{nn({\mathcal{X}})}=p_{nn({\mathcal{X}}),K+1},$ (24c)
where $\bm{L}^{{{w}}_{k}}:{\mathbb{R}}^{D\cdot n_{k}}\rightarrow{\mathbb{R}}^{D\cdot n_{k}}$ is the stacking of $D$ times $L^{k}$, i.e.,
$\bm{L}^{{{w}}_{k}}\left(({{z}}_{1}^{T},\ldots,{{z}}_{D}^{T})^{T})=(L^{{{w}}_{k}}({{z}}_{1})^{T},\allowbreak\ldots,L^{{{w}}_{k}}({{z}}_{D})^{T}\right)^{T}.$
@@
That is, $p_{nn({\mathcal{X}})}$ is computed by stacking $D$ times the neural network for the inputs in $\mathcal{X}$}. Following the same steps as in the previous section, $p_{nn({\mathcal{X}})}$ can then be approximated by the Gaussian Mixture distribution $q_{nn({\mathcal{X}})}$, for $k\in\{1,\ldots,K\}$ defined as
@@
$q_{nn({\mathcal{X}}),1}={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{\mathrm{vec}}({\mathcal{X}})}}\left[\bm{L}^{{{w}}_{0}}({{{z}}}_{0})\#p_{{{w}}_{0}}\right],$ (Initialization and Affine operation) (25a)
$d_{nn({\mathcal{X}}),k}=\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}\#{G}_{{M}}(q_{nn({\mathcal{X}}),k}),$ (Compression and Signature operations) (25b)
@@
$q_{nn({\mathcal{X}}),k+1}={\mathbb{E}}_{{{{z}}}_{k}\sim d_{nn({\mathcal{X}}),k}}\left[\bm{L}^{{{w}}_{k}}({\sigma}({{{z}}}_{k}))\#p_{{{w}}_{k}}\right],$ (Activation and Affine operations) (25c)
$q_{nn({\mathcal{X}})}=q_{nn({\mathcal{X}}),K+1},$ (Output) (25d)
@@
where $\bm{{\mathcal{C}}}_{k}=\{{\mathcal{C}}_{k,i}\subset{\mathbb{R}}^{D\cdot n_{k}}\}_{i=1}^{M_{k}}$ are the sets of signature locations at layer $k$, and $M_{k}$$\in\mathbb{N}$ is the compression size. Theorem 17 implies the following corollary, which bounds the approximation error of $q_{nn({\mathcal{X}})}$.
Corollary 20 (of Theorem 17)
Let $p_{nn({\mathcal{X}})}$ be the joint distribution of a SNN with $K$ hidden layers over a finite set of input points ${\mathcal{X}}=\{{{x}}_{i}\}_{i=1}^{D}\subset{\mathbb{R}}^{n_{0}}$ and $q_{nn({\mathcal{X}})}$ be a Gaussian mixture distribution built according to Eqn (25). Iteratively define $\hat{{\mathbb{W}}}_{2}$ as
$\hat{{\mathbb{W}}}_{2,1}=0,$ (26a)
$\hat{{\mathbb{W}}}_{2,k+1}={\mathcal{L}}_{k}\left[\hat{{\mathbb{W}}}_{2,k}+{\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}}),k},{G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)\right$.
$\hskip 142.26378pt\left.+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)\right],\qquad\forall k\in\{1,\ldots,K\},$ (26b)
$\hat{{\mathbb{W}}}_{2}=\hat{{\mathbb{W}}}_{2,K+1},$ (26c)
where $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}}}$ is defined in Eqn (14), ${\mathbb{MW}}_{2}$ as defined in Eqn (2), and ${\mathcal{L}}_{k}$ is as defined Eqn (22). Then, it holds that
Note that the effect of multiple input points on the error introduced by $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$ and ${\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}}),k},{G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$ depends on the specific choice of points in $\mathcal{X}$}. For points in $\mathcal{X}$} where $q_{nn({\mathcal{X}}),k}$ is highly correlated, the error will be similar to that of a single-input case. However, if there is little correlation, the error bound will increase linearly with the number of points.
6.4 Convergence Analysis
In Theorem 17 and Corollary 20, we derived error bounds on the distance between $q_{nn({\mathcal{X}})}$ and $p_{nn({\mathcal{X}})}$, which depend on the signature locations and on the size of the compressed mixtures. The following theorem shows that by appropriately selecting these quantities, the error bound can be made arbitrarily small.
Theorem 21
Let $p_{nn({\mathcal{X}})}$ be the output distribution of a SNN with $K$ hidden layers for a finite set of inputs ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$. For any $\epsilon>0$, for layer $k\in\{1,\ldots,K\}$, choose a compression size $M_{k}\in\mathbb{N}$ such that
$\epsilon_{k}\coloneqq\frac{\epsilon}{K\prod_{i=k}^{K}{\mathcal{L}}_{i}}-{\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}}),k},{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\right)>0,$ (27)
where ${\mathcal{L}}_{i}$ is defined in Eqn (22). Choose per-component, per-dimension signature grid sizes $N_{k,i,j}\in\mathbb{N}$ satisfying
$\sum_{i=1}^{M_{k}}{\tilde{\pi_{k}}^{{(i)}}}\sum_{j=1}^{n_{k}}{{{\lambda}}_{k,i}^{{(j)}}}\left(\frac{4\log N_{k,i,j}}{N_{k,i,j}^{2}}+r(N_{k,i,j})\right)\leq\epsilon_{k}^{2},$ (28)
where $\tilde{\pi}_{k}$ are the mixture weights, ${{\lambda}}_{k,i}$ is the vector of eigenvalues of the covariance matrix of the $i$-th component of $q_{nn({\mathcal{X}}),k}$, and $r(\cdot)$ is given in Eqn (17). Construct for each component $i$ of $q_{nn({\mathcal{X}}),k}$ a set of signature locations ${\mathcal{C}}_{k,i}\subset{\mathbb{R}}^{n_{k}}$ as in Eqn (12), define $\bm{{\mathcal{C}}}_{k}=\{{\mathcal{C}}_{k,i}\}_{i=1}^{M_{k}},$ and let $q_{nn({\mathcal{X}}),k}$ be as in Eqn (25). Then, for $\hat{{\mathbb{W}}}_{2}$ defined in Eqn (26), it holds that
${\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}}),K+1}\right)\leq\hat{\mathbb{W}}_{2}\leq\epsilon.$ (29)
To prove Theorem 21, we use that Corollary 20 implies the layer-wise decomposition
$\hat{\mathbb{W}}_{2}=\sum_{k=1}^{K}\Bigg{(}\prod_{i=k}^{K}{\mathcal{L}}_{i}\Bigg{)}\Big{[}{\mathbb{MW}}_{2}\big{(}q_{nn({\mathcal{X}}),k},{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\big{)}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\big{(}{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\big{)}\Big{]}.$
That is, $\hat{\mathbb{W}}_{2}$ equals the sum of compression and signature errors at layer $k$, scaled by the product of Lipschitz constants of the remaining layers $\prod_{i=k}^{K}{\mathcal{L}}_{i}$. If at each layer $k$ the compression size $M_{k}$ and grid sizes $N_{k,i,j}$ are such that
${\mathbb{MW}}_{2}\big{(}q_{nn({\mathcal{X}}),k},{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\big{)}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\big{(}{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\big{)}\leq\frac{\epsilon}{K\prod_{i=k}^{K}{\mathcal{L}}_{i}},$
it follows that $\hat{\mathbb{W}}_{2}\leq\epsilon$.
For any choice of $M_{k}$ such that $\epsilon_{k}>0$ (Eqn (27)), Corollary 15 allows us to select grid sizes $N_{k,i,j}$ so that $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\big{(}{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\big{)}\leq\epsilon_{k}$, thereby satisfying the per-layer requirement and yielding $\hat{\mathbb{W}}_{2}\leq\epsilon$. Applying no compression (i.e., setting $M_{k}$ equal to the original number of mixture components in $q_{nn({\mathcal{X}}),k}$) ensures $\epsilon_{k}>0$ but can cause exponential growth in the total number of signature locations across depth, since signatures are taken component-wise. In practice, we therefore compress only redundant or low-weight components, i.e., retaining distinct modes, which preserves most of the signature budget while substantially reducing the number of grids (see Section 8.1).
Remark 22
The number of signature locations required at layer $k$ in Theorem 21 is essentially inversely proportional to the layer accuracy parameter $\epsilon_{k}$, and hence to the target accuracy $\epsilon$ (up to the distribution of the budget across layers). [^14] At layer $k$, the signature contribution behaves like $O\left(\tfrac{\log N_{k,i,j}}{N_{k,i,j}^{2}}\right)$ for large $N_{k,i,j}$. Solving $\log N_{k,i,j}/N_{k,i,j}^{2}\approx\epsilon_{k}^{2}$ gives $N_{k,i,j}=O\big{(}\epsilon_{k}^{-1}\sqrt{\log(1/\epsilon_{k})}\big{)}$: a $1/\epsilon_{k}$ rate with only a mild $\sqrt{\log}$ factor. In the worst-case scenario where each of the $n_{k}$ dimensions of a component requires the same grid size, the total number of signature locations for that component scales as $\prod_{j=1}^{n_{k}}N_{k,i,j}=O\big{(}\epsilon_{k}^{-n_{k}}(\log(1/\epsilon_{k}))^{n_{k}/2}\big{)}$. In practice, this curse of dimensionality is strongly mitigated because Condition (28) weights each dimension by its eigenvalue. Empirically, most variance concentrates in a few neurons per layer, so the effective dimensionality is lower and many low-variance dimensions require only a single point ($N_{k,i,j}=1$). Additionally, non-uniform and adaptive (e.g., variance-proportional) grids further reduce required sizes while preserving ${\mathbb{W}}_{2}$ accuracy (see Remark 12). In Section 8.1, we empirically analyze how this nominal exponential dependence is mitigated in practice.
input : $p_{nn({\mathcal{X}})}$
output : $q_{nn({\mathcal{X}})}$ and $\hat{{\mathbb{W}}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)$
begin
 1 Initialize $q_{nn({\mathcal{X}}),1}$ as in Eqn (25a) and $\hat{{\mathbb{W}}}_{2,1}$ as in Eqn (21a)
 for $k\in\{1,\ldots,K\}$ do
    2 Construct ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$ using compression Algorithm 3
    3 Compute ${\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}}),k},{G}_{M_{k}}(q_{nn({\mathcal{X}}),k})\right)$ as in Eqn (2)
    4 Construct signature locations $\bm{{\mathcal{C}}}_{k}$ and the signature $d_{nn({\mathcal{X}}),k}$ of ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$, and compute the signature error $\hat{{\mathbb{W}}}_{2,{\Delta}_{\bm{{\mathcal{C}}}_{k}}}({G}_{M_{k}}(q_{nn({\mathcal{X}}),k}))$ all via Algorithm 2
    5 Update $q_{nn({\mathcal{X}}),k+1}$ as in Eqn (25c)
    6 Compute ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{\frac{1}{2}}$ as described in Appendix C
    7 Update $\hat{{\mathbb{W}}}_{2,k+1}$ as in Eqn (21b)
 8Return $q_{nn({\mathcal{X}})}=q_{nn({\mathcal{X}}),K+1}$ and $\hat{{\mathbb{W}}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)=\hat{{\mathbb{W}}}_{2,K+1}$
Algorithm 1 Gaussian Mixture Approximation of a SNN
7 Algorithmic Framework
In this section, the theoretical methodology to solve Problem 1 developed in the previous sections is translated into an algorithmic framework. First, following Eqn (25), we present an algorithm to construct the Gaussian mixture approximation of a SNN with error bounds in the 2-Wasserstein distance, including details on the compression step. Then, we rely on the fact that the error bound resulting from our approach is piecewise differentiable with respect to the parameters of the SNN to address the complementary problem highlighted in Remark 4. This problem involves deriving an algorithm to optimize the parameters of a SNN such that the SNN approximates a given GMM.
7.1 Construction of the GMM Approximation of a SNN
The procedure to build a Gaussian mixture approximation $q_{nn({\mathcal{X}})}$ of $p_{nn({\mathcal{X}})}$, i.e. the output distribution of a SNN at a set of input points $\mathcal{X}$} as in Eqn (25), is summarized in Algorithm 1. The algorithm consists of a forward pass over the layers of the SNN. First, the Gaussian output distribution after the first affine layer is constructed (line 1). Then, for each layer $k>0$, $q_{nn({\mathcal{X}}),k}$ is compressed to a mixture of size $M$ (line 2); the signature of ${G}_{{M_{k}}}(q_{nn({\mathcal{X}})})$ is computed (line 4); and the signature is passed through the activation and affine layer to obtain $q_{nn({\mathcal{X}}),k+1}$ (line 5). Parallel to this, error bounds on the 2-Wasserstein distance are computed for the compression operation (line 3), signature operation (line 4), and affine operation (line 6), which are then composed into a bound on ${\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}}),k+1},q_{nn({\mathcal{X}}),k+1}\right)$ that is propagated to the next layer (line 7).
Signature Operation
The signature operation on $q_{nn({\mathcal{X}}),k}$ in line 4 of Algorithm 1 is performed according to the steps in Algorithm 2. That is, the signature of $q_{nn({\mathcal{X}}),k}$ is taken as the mixture of the signatures of the components of $q_{nn({\mathcal{X}}),k}$ (lines 1-6). For each Gaussian component of the mixture, to ensure an analytically exact solution for the Wasserstein distance resulting from the signature, we position its signature locations on a grid that aligns with the covariance matrix’s basis vectors, i.e., the locations satisfy the constraint as in Eqn (12) of Corollary 11. In particular, the grid of signatures is constructed so that the Wasserstein distance from the signature operation is minimized. According to the following proposition, the problem of finding this optimal grid can be solved in the transformed space, in which the Gaussian is independent over its dimensions.
@@ 7.1 Construction of the GMM Approximation of a SNN
For $\mathcal{N}({m},\Sigma)\in{\mathcal{P}}({\mathbb{R}}^{n})$, let matrix ${{T}}\in{\mathbb{R}}^{n\times n}$ be such that ${{T}}=\mathrm{diag}({{\lambda}})^{-\frac{1}{2}}{{V}}^{T}$ where $\mathrm{diag}({{\lambda}})={{V}}^{T}\Sigma{{V}}$ is a diagonal matrix whose entries are the eigenvalues of $\Sigma$, and V{{V}} is the corresponding orthogonal (eigenvector) matrix. Then,
@@
$\underset{\begin{subarray}{c}\{{\mathcal{C}}^{1},\ldots,{\mathcal{C}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|{\mathcal{C}}^{i}|=N\end{subarray}}{\mathrm{argmin}}{\mathbb{W}}_{2}\left({\Delta}_{\mathrm{Post}({\mathcal{C}}^{1}\times\ldots\times{\mathcal{C}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma)\right)=\left\{{\mathcal{C}}^{*}(N_{1}^{*}),\ldots,{\mathcal{C}}^{*}(N_{n}^{*})\right\},$